complete visual alsoProblem Statement A 5.0-4C point charge is placed at the 0.00 cm mark of a meter stick and a -4.0-ųC point charge is placed a
Answer #1

Solution :

Visual representation :

.

.

Here we have :

q1 = 5 μC

q2 = - 4 μC

Let the net electric field at point P is zero.

Then : E1 = E2

\therefore \ \frac{kq_1}{x^2}=\frac{kq_2}{(x-0.50)^2}

\therefore \ \frac{(5\ \mu C)}{x^2}=\frac{(4\ \mu C)}{(x-0.50)^2}

\therefore \ (5)(x-0.50)^2=(4)x^2

∴ (5)(x2 - x + 0.25) = 4 x2

∴ 5 x2 - 5x + 1.25 = 4 x2

∴ x2 - 5x + 1.25 = 0

Solving this quadratic equation for x gives : x = 4.736 m

Therefore, Electric field will be Zero at 4.736 m or 473.6 cm